# RETENTION: Resource-Efficient Tree-Based Ensemble Model Acceleration with Content-Addressable Memory

### v1

@@ Abstract
Although deep learning has demonstrated remarkable capabilities in learning from unstructured data, modern tree-based ensemble models remain superior in extracting relevant information and learning from structured datasets. While several efforts have been made to accelerate tree-based models, the inherent characteristics of the models pose significant challenges for conventional accelerators. Recent research leveraging content-addressable memory (CAM) offers a promising solution for accelerating tree-based models, yet existing designs suffer from excessive memory consumption and low utilization. This work addresses these challenges by introducing RETENTION, an end-to-end framework that significantly reduces CAM capacity requirement for tree-based model inference. We propose an iterative pruning algorithm with a novel pruning criterion tailored for bagging-based models (e.g., Random Forest), which minimizes model complexity while ensuring controlled accuracy degradation. Additionally, we present a tree mapping scheme that incorporates two innovative data placement strategies to alleviate the memory redundancy caused by the widespread use of don’t care states in CAM. Experimental results show that implementing the tree mapping scheme alone achieves $1.46\times$ to $21.30\times$ better space efficiency, while the full RETENTION framework yields $4.35\times$ to $207.12\times$ improvement with less than 3% accuracy loss. These results demonstrate that RETENTION is highly effective in reducing CAM capacity requirement, providing a resource-efficient direction for tree-based model acceleration.
@@ I Introduction
Structured (i.e., tabular) data is one of the most prevalent formats in data science. It is typically represented as a matrix, where each row corresponds to an instance and all instances share the same set of features across columns. This format is widely used in fields such as finance, medicine, and scientific research, as its structured nature enables efficient processing. For example, sensors generate data in tabular format to support real-time analysis for AI-driven closed-loop control. Despite significant advancements in deep learning for unstructured data (e.g., text, images, and speech), studies [1, 2] have shown that modern tree-based ensemble models consistently outperform deep learning in several tasks. In fact, tree-based models remain the state-of-the-art for classification and regression tasks involving medium-sized structured datasets [2], and are often favored over deep learning due to their high interpretability [3], particularly in sensitive applications where understanding model decisions is critical [4]. Tree-based models are widely applied in domains such as scientific research [5], credit card fraud detection [6], and machinery fault diagnosis [7]. A recent survey [8] found that over 74% of data scientists prefer tree-based models, whereas fewer than 40% opt for neural networks, underscoring their continued importance.
@@
Recently, there has been significant interest in developing emerging non-volatile-memory-based (NVM-based) IMC accelerators [15, 16, 17], as the characteristics of emerging NVM (e.g., high density, low power consumption, and low cost) make it well-suited for resource-constrained environments. Among potential candidates, non-volatile ternary content-addressable memory (nvTCAM) [18, 19] exhibits the best fit for accelerating tree-based models due to its capability to perform sequence matching with extremely high parallelism and its high reliability. Since every root-to-leaf path within tree-based models can be encoded into a binary sequence, nvTCAM can traverse all the paths in one shot, enabling unprecedented acceleration for model inference. Nevertheless, to support in-memory search, data must be organized in a specialized format (refer to Section II-B for further explanation). Although nvTCAM effectively accelerates tree-based model inference in a cost- and energy-efficient manner, a considerable portion of memory cells is allocated for format alignment rather than storing actual model data. This leads to excessive memory consumption with substantial redundancy, which is highly inefficient and requires further optimization.
To achieve resource-efficient acceleration, the enormous CAM capacity requirement with considerable redundancy is the major obstacle that must be addressed before practical implementation. Since model complexity is highly correlated with memory consumption, pruning models can significantly mitigate this issue. Modern tree-based models often employ pruning algorithms, such as limiting maximum depth or minimum impurity decrease [20], in order to reduce overfitting and simplify model structure. While these techniques are well-suited for boosting-based ensemble models (e.g., XGBoost), where each tree in the ensemble is trained to correct the errors of the previous ones, applying them to bagging-based tree ensemble models (e.g., Random Forest [21]) can lead to severe accuracy degradation due to the independent training of each tree. On the other hand, memory redundancy arises from structuring data for in-memory search, suggesting that a tailored data placement strategy could further reduce CAM capacity requirement. While several studies [22, 23, 24, 25, 26] have explored accelerating tree-based models using non-volatile CAM, little attention has been paid to the issue of memory redundancy, and existing data placement strategies fail to address this challenge effectively and efficiently. Mitigating redundancy requires a solution that not only reduces model complexity but also optimizes memory utilization, paving the way for a more resource-efficient acceleration.
@@
To the best of our knowledge, this is the first work that systematically optimizes data placement strategies for tree-based model acceleration with CAM, explicitly considering the tradeoffs between memory redundancy and processing overhead. RETENTION is evaluated on Random Forest and XGBoost using five datasets. Although validated with these two models, the framework can be generalized to other ensemble models with similar structures such as LightGBM [28] and CatBoost [29]. Experimental results show that our tree mapping scheme alone achieves $1.46\times$ to $21.30\times$ better space efficiency, while the complete framework with pruning and tree mapping optimization improves space efficiency by $4.35\times$ to $207.12\times$ with less than 3% accuracy degradation. These results demonstrate that RETENTION effectively reduces CAM capacity requirement, making tree-based model acceleration with nvTCAM more resource-efficient and feasible for resource-constrained environments.
@@ II-A Decision Tree, Bagging, and Boosting
Figure 1: Visualization of decision tree inference acceleration with CAM.
@@
Bagging (i.e., bootstrap aggregation) [31] is a widely used algorithm for tree-based ensemble model training. It trains multiple fully-grown decision trees independently, each on a unique subset of the dataset generated through bootstrapping (i.e., sampling with replacement). When making a prediction, each tree produces an individual result, and the ensemble determines the final output via majority voting (if classification) or averaging (if regression). By training trees on different subsets of data, bagging enhances robustness against overfitting, reducing the need for pruning. When additional model compression is required, early-stopping (i.e., pre-pruning) is preferred to pruning constructed trees (i.e., post-pruning), as conventional post-pruning algorithms ignore the collective behavior of the model and can erode the diversity that bagging-based model relies on. Additionally, bagging enables OOB estimation, an inherent validation method that eliminates the need for a separate validation set. Since each tree is trained on a subset of data, the unused instances can serve as the validation set, which is later passed through the corresponding tree for evaluation. By aggregating predictions from all trees that did not train on a given instance, OOB estimation provides an unbiased measure of model performance, making it particularly useful when data is limited. Bagging alleviates overfitting, while the ensemble compensates for the missing information in individual trees, resulting in improved accuracy compared to a single decision tree. Additionally, since each tree in a bagging-based model is trained independently and contributes equally to the final prediction, models such as Random Forest perform exceptionally well on noisy and small datasets. However, they struggle to capture complex patterns, as they do not fully exploit interactions among trees within the ensemble. Moreover, the averaging mechanism limits the model performance in regression tasks, highlighting the need for an alternative.
@@
Another class of tree ensemble model primitives is the boosting-based model, which is commonly utilized for both classification and regression tasks. In contrast to bagging, which trains multiple fully-grown decision trees independently, boosting [32] trains shallow decision trees sequentially, with each tree attempting to correct the errors of its predecessors. During training, boosting fits the first tree on the entire dataset, and each subsequent tree is trained to correct the errors of its predecessors, with errors being reevaluated at each step as new trees are added to the ensemble. When making a prediction, the final result is computed as a weighted sum of individual tree predictions rather than relying on majority voting or averaging, as trees in boosting-based models contribute with different importance weights. While the error correction mechanism and the sequential training process adapt the models to task complexity, they are more susceptible to overfitting and therefore rely heavily on model pruning. However, since the training phase leverages interactions among trees, post-pruning can disrupt these inter-tree dependencies, leading to catastrophic accuracy degradation. Consequently, boosting-based models often opt for pre-pruning algorithms such as limiting maximum depth and minimum impurity decrease. However, modern boosting implementations, such as XGBoost, can easily contain thousands of trees, resulting in substantial memory consumption. Additionally, the irregular and non-deterministic memory access patterns in tree traversal make efficient execution on conventional hardware challenging. To mitigate these inefficiencies, content-addressable memory (CAM) enables massively parallel in-memory search, significantly improving inference speed and efficiency.
@@ II-B Ternary CAM and Analog CAM
Content-addressable memory (CAM) [14, 33, 34] is a specialized memory architecture widely used in various applications. Unlike conventional memory, which retrieves data based on a given address, CAM allows simultaneous comparison of an input query sequence against all stored binary sequences, and returns the addresses or associated data of matching entries. Data in CAM is stored row-wise, with comparisons performed in parallel on a column basis. This high degree of parallelism makes CAM particularly well-suited for applications requiring rapid lookups, such as network routing, database indexing, and cache systems.
Ternary content-addressable memory (TCAM) [35] extends the functionality of binary CAM by introducing a third state, don’t care, denoted as X. Bits set to the X state are treated as matches during comparison, enabling greater flexibility in search operations. Since each root-to-leaf path in a tree-based model represents a sequence of conditions an input must satisfy to reach the corresponding prediction, TCAM can efficiently traverse these paths, making it a compelling solution for tree-based model acceleration. Prior work [22] proposed mapping each path to a TCAM row, with each column representing a unique condition. Unencountered conditions within a path are assigned the X state as they do not affect the traversal process. When an input instance is received, it is first encoded into a binary sequence that aligns with the condition order stored in TCAM. This encoded sequence is then fed into TCAM for in-memory search. The middle part of Fig. 1 illustrates how TCAM serves as a decision tree accelerator. The decision tree at the left part of Fig. 1 is mapped to the TCAM, where four unique conditions are assigned to separate columns, and each path is represented as a row. Conditions are encoded as 0, 1, and X, corresponding to False, True, and don’t care, respectively. Once an input is received, the features are encoded into a binary sequence representing the condition check results, and each bit is then delivered to the corresponding column to perform in-memory search. Since all bits in the second row are matched, $Path_{2}$ is the traversal path for the input instance.
@@
Analog content-addressable memory (ACAM) [41] follows a similar structural design to TCAM, but differs in its data representation. Instead of discrete states (0, 1, X), each ACAM cell can store a range of analog values. A cell is considered matched if the given input value falls within the stored range, meaning that a min-to-max range in ACAM functions analogously to the X state in TCAM. Previous research [23] has demonstrated the potential of ACAM for accelerating Random Forest. Similar to [22], each root-to-leaf path is mapped to an ACAM row. However, unlike TCAM, where each column represents a feature-threshold condition, ACAM columns directly correspond to features. When an input instance is received, a digital-to-analog converter first transforms the features into analog voltages. ACAM then performs in-memory search operations to retrieve the final result. The right part of Fig. 1 provides an overview of ACAM serving as the accelerator for the decision tree in the left part of Fig. 1. In contrast to TCAM, ACAM only requires three columns to represent the features, and each of the cells in ACAM stores an analog range instead.
However, despite its space efficiency, ACAM requires additional peripheral circuits, and its larger cell size does not necessarily offer better area efficiency than TCAM. Furthermore, analog computing is susceptible to various non-idealities, such as stuck-at faults, IR drop, thermal noise, shot noise, and random telegraph noise [42]. While mitigation techniques exist [43], they often introduce additional memory overhead and offer limited effectiveness. Due to these reliability challenges, ACAM is not yet viable for real-world implementation. Therefore, this work opts for TCAM as the accelerator. Still, RETENTION remains compatible with ACAM after adapting the condition-based operations to operate on features only.
@@ II-C Observation
To accelerate tree-based model inference using TCAM, each root-to-leaf path is mapped to a TCAM row, with each column representing a unique condition within the ensemble. Since a single path encounters only a minimal fraction of the conditions within the ensemble, the majority of TCAM cells store the X state for format alignment, leading to significant redundancy. In fact, mapping a tree-based model to TCAM without optimizations requires at least $\textit{paths}\times\#\textit{unique\_conditions}$ bits, far exceeding the memory capacity available in resource-constrained environments. As shown in Fig. 1, mapping a decision tree to a CAM with perfect capacity leads to over one-third of cells storing the X state. Moreover, due to the extremely high write latency of nvTCAM, real-time writing severely degrades performance, making it critical to fit the entire model into nvTCAM. Table I presents an experiment that highlights the substantial memory requirement and the overwhelming redundancy caused by the widespread use of the X states in TCAM. In this experiment, Random Forest models with 100 decision trees require up to 700MB of memory when mapped with naive unified mapping (refer to Section III-D1 for further explanation), and all exhibited redundancy of more than 96%. These results highlight that directly deploying nvTCAM as a tree-based model accelerator is both infeasible and inefficient in resource-constrained environments.
Previous work DT2CAM [22] introduced a framework for mapping decision trees to TCAM. However, instead of addressing memory redundancy, it only proposed an energy-saving precharge mechanism that disables unmatched rows to conserve power. Pedretti et al. [23] proposed mapping Random Forests to ACAM, incorporating techniques such as limiting tree depth and reordering features while eliminating empty rows to mitigate memory redundancy. However, pre-pruning algorithms can result in severe accuracy degradation when implementing on bagging-based models (refer to Section III-B for further explanation), while alternative pruning strategies could achieve better accuracy within the same memory budget. Furthermore, although feature reordering groups redundant cells and empty-row elimination reduces capacity requirement, the associated computational overhead for result retrieval is relatively high. In fact, comparable computational overhead could yield even greater memory savings. A follow-up study by Pedretti et al. [24] proposed assigning individual trees to separate ACAMs, improving space efficiency. While this approach effectively reduces redundancy across different trees, redundancy within each separate tree remains, leaving room for further optimization. Addressing this residual redundancy could lead to more reductions in memory consumption while maintaining computational efficiency.

### v2

@@ Abstract
Although deep learning has demonstrated remarkable capability in learning from unstructured data, modern tree-based ensemble models remain superior in extracting relevant information and learning from structured datasets. While several efforts have been made to accelerate tree-based models, the inherent characteristics of the models pose significant challenges for conventional accelerators. Recent research leveraging content-addressable memory (CAM) offers a promising solution for accelerating tree-based models, yet existing designs suffer from excessive memory consumption and low utilization. This work addresses these challenges by introducing RETENTION, an end-to-end framework that significantly reduces CAM capacity requirement for tree-based model inference. We propose an iterative pruning algorithm with a novel pruning criterion tailored for bagging-based models (e.g., Random Forest), which minimizes model complexity while ensuring controlled accuracy degradation. Additionally, we present a tree mapping scheme that incorporates two innovative data placement strategies to alleviate the memory redundancy caused by the widespread use of don’t care states in CAM. Experimental results show that implementing the tree mapping scheme alone reduces CAM capacity requirement by $1.46\times$ to $21.30\times$, while the full RETENTION framework achieves $4.35\times$ to $207.12\times$ reduction with less than 3% accuracy loss. These results demonstrate that RETENTION is highly effective in minimizing CAM resource demand, providing a resource-efficient direction for tree-based model acceleration.
@@ I Introduction
Structured (i.e., tabular) data is one of the most prevalent formats in data science. It is typically represented as a matrix, where each row corresponds to an instance and all instances share the same set of features across columns. This format is widely used in fields such as finance, medicine, and scientific research, as its structured nature enables efficient processing. For example, sensors generate data in tabular format to support real-time analysis for AI-driven closed-loop control. Despite significant advancements in deep learning for unstructured data (e.g., text, images, and speech), studies [39, 19] have shown that modern tree-based ensemble models consistently outperform deep learning in several tasks. In fact, tree-based models remain the state-of-the-art for classification and regression tasks involving medium-sized structured datasets [19], and are often favored over deep learning due to their high interpretability [28], particularly in sensitive applications where understanding model decisions are critical [17]. Tree-based models are widely applied in domains such as scientific research [50], credit card fraud detection [1], machinery fault diagnosis [48], etc. A recent survey [21] found that over 74% of data scientists prefer tree-based models, whereas fewer than 40% opt for neural networks, underscoring their continued importance.
@@
Recently, there has been significant interest in developing emerging non-volatile-memory-based (NVM-based) IMC accelerators [38, 14, 26], as the characteristics of emerging NVM (e.g., high density, low power consumption, and low cost) make it well-suited for resource-constrained environments. Among potential candidates, non-volatile ternary content-addressable memory (nvTCAM) [49, 51, 11] exhibits the best fit for accelerating tree-based models due to its capability to perform sequence matching with extremely high parallelism, energy efficiency, and reliability. Since every root-to-leaf path within tree-based models can be encoded into a binary sequence, nvTCAM can traverse all the paths in one shot, enabling unprecedented acceleration for model inference. Nevertheless, to support in-memory search, data must be organized in a specialized format (refer to Section II-B for detailed explanation). Although nvTCAM effectively accelerates tree-based model inference in a cost- and energy-efficient manner, a considerable portion of memory cells is allocated for format alignment rather than storing actual model data. This leads to excessive memory consumption with substantial redundancy, which is highly inefficient and requires further optimization.
To achieve resource-efficient acceleration, the enormous CAM capacity requirement with considerable redundancy is the major obstacle that must be addressed before practical implementation. Since model complexity is highly correlated with memory consumption, pruning models can significantly mitigate this issue. Modern tree-based models often employ pruning algorithms, such as limiting maximum depth or minimum impurity decrease [10], in order to reduce overfitting and simplify model structure. While these techniques are well-suited for boosting-based ensemble models (e.g., XGBoost), where each tree in the ensemble is trained to correct the errors of the previous ones, applying them to bagging-based ensemble models (e.g., Random Forest [9]) can lead to severe accuracy degradation due to the independent training of each tree (refer to Section III-B for detailed explanation). On the other hand, memory redundancy arises from structuring data for in-memory search, suggesting that a tailored data placement strategy could further reduce CAM capacity requirement. While several studies [35, 31, 32, 41, 46] have explored accelerating tree-based models using non-volatile CAM, little attention has been paid to the issue of memory redundancy, and existing data placement strategies fail to address this challenge effectively and efficiently. Mitigating redundancy requires a solution that not only reduces model complexity but also optimizes memory utilization, paving the way for a more resource-efficient acceleration.
@@
To the best of our knowledge, this work is the first to explicitly consider the tradeoffs between memory redundancy and processing overhead, offering data placement strategies for different optimization criteria. RETENTION is evaluated on Random Forest and XGBoost using five datasets. Although validated with these two models, the framework can be generalized to other ensemble models with similar structures such as LightGBM [24] and CatBoost [33]. Experimental results show that our tree mapping scheme alone reduces CAM capacity requirement by $1.46\times$ to $21.30\times$, while the full RETENTION framework with model pruning and mapping optimization achieves $4.35\times$ to $207.12\times$ reduction with less than 3% accuracy degradation. These results demonstrate that RETENTION effectively reduces CAM capacity requirement, making tree-based model acceleration with CAM more resource-efficient and feasible for resource-constrained environments.
@@ II-A Decision Tree, Bagging, and Boosting
Figure 1: Visualization of decision tree inference acceleration with TCAM and ACAM.
@@
Bagging (i.e., bootstrap aggregation) [7] is a widely used algorithm for tree-based ensemble model training. It trains multiple fully-grown decision trees independently, each on a unique subset of the training set generated through bootstrapping (i.e., sampling with replacement). During inference, each tree produces an individual result, and the ensemble determines the final output via majority voting (if classification) or averaging (if regression). By training trees on different subset, bagging enhances robustness against overfitting, reducing the need for pruning. When additional model compression is required, early-stopping (i.e., pre-pruning) is preferred to pruning constructed trees (i.e., post-pruning), as conventional post-pruning algorithms ignore the collective behavior of the model and can erode the diversity that bagging-based model relies on. Additionally, bagging enables OOB estimation, an inherent validation method that eliminates the need for a separate validation set. Since each tree is trained on a subset of data, the unused instances can serve as the validation set, which is later passed through the corresponding tree for evaluation. By aggregating predictions from all trees that did not train on a given instance, OOB estimation provides an unbiased measure of model performance, making it particularly advantageous in data-constrained scenarios. Bagging alleviates overfitting, while the ensemble compensates for the missing information in individual trees, resulting in improved accuracy compared to a single decision tree. Additionally, since each tree in a bagging-based model is trained independently and contributes equally to the final prediction, models such as Random Forest perform exceptionally well on noisy and small datasets. However, they struggle to capture complex patterns, as they do not fully exploit interactions among trees within the ensemble. Moreover, the averaging mechanism limits model performance in regression tasks, highlighting the need for an alternative.
@@
Another class of tree-based ensemble model primitives is the boosting-based model, which is commonly utilized for both classification and regression tasks. In contrast to bagging, which trains multiple fully-grown decision trees independently, boosting [36] trains shallow decision trees sequentially, with each tree attempting to correct the errors of its predecessors. During training, boosting fits the first tree on the entire dataset, and each subsequent tree is trained to correct the errors of the previous ones, with errors being reevaluated at each step as new trees are added to the ensemble. When making a prediction, the final result is computed as a weighted sum of individual tree predictions rather than relying on majority voting or averaging, as trees in boosting-based models contribute with different importance weights. While the error correction mechanism and the sequential training process adapt the models to task complexity, they are more susceptible to overfitting and therefore rely heavily on model pruning. However, since the training phase leverages interactions among trees, post-pruning can disrupt these inter-tree dependencies, leading to catastrophic accuracy degradation. Consequently, boosting-based models often opt for pre-pruning algorithms such as limiting maximum depth and minimum impurity decrease. Nevertheless, modern boosting implementations, such as XGBoost, can easily contain thousands of trees, resulting in substantial memory consumption. Additionally, the irregular tree structures and non-deterministic memory access patterns make efficient inference on conventional hardware particularly challenging. These limitations underscore the potential of CAM, which replaces the time- and energy-consuming memory accesses with efficient in-memory search.
@@ II-B Ternary CAM and Analog CAM
Content-addressable memory (CAM) [30, 23, 29] is a specialized memory architecture widely used in various applications. Unlike conventional memory, which retrieves data based on a given address, CAM allows simultaneous comparison of an input query sequence against all stored sequences, and returns the addresses or associated data of matching entries. Data in CAM is stored row-wise, with comparisons performed in parallel on a column basis. This high degree of parallelism makes CAM particularly well-suited for applications requiring rapid lookups, such as network routing, database indexing, and cache systems.
Ternary content-addressable memory (TCAM) [3] extends the functionality of binary CAM by introducing a third state, don’t care, denoted as X. Bits set to the X state are treated as matches during comparison, enabling greater flexibility in search operations. Since each root-to-leaf path in a tree-based model represents a sequence of conditions an input must satisfy to yield the corresponding prediction, TCAM can efficiently traverse these paths, making it a compelling solution for tree-based model acceleration. Prior work [35] proposed mapping each path to a TCAM row, with each column representing a unique condition. Unencountered conditions of a path are assigned the X state as they do not affect the traversal process. When an input instance is received, it is first encoded into a binary sequence representing the condition check results. This sequence is then used to construct TCAM input queries, which are dispatched to the respective TCAMs for in-memory search. The middle part of Fig. 1 illustrates how TCAM serves as a decision tree accelerator. The decision tree at the left part of Fig. 1 is mapped to the TCAM, where four unique conditions are assigned to separate columns, and each path is represented as a row. Conditions are encoded as 0, 1, and X, corresponding to False, True, and don’t care, respectively. Once an input is received, the features are encoded into a binary sequence representing the condition check results, and each bit is then delivered to the corresponding column to perform in-memory search. Since all cells in the second row are matched, $Path_{2}$ is the traversal path for the input instance.
@@
Analog content-addressable memory (ACAM) [27] follows a similar structural design to TCAM, but differs in its data representation. Instead of discrete states (0, 1, X), each ACAM cell can store a range of analog values. A cell is considered matched if the given input value falls within the stored range, meaning that a min-to-max range in ACAM functions analogously to the X state in TCAM. Previous research [31] has demonstrated the potential of ACAM for accelerating Random Forest. Similar to [35], each root-to-leaf path is mapped to an ACAM row. However, unlike TCAM, where each column represents a feature-threshold condition, ACAM columns directly correspond to features. When an input instance is received, a digital-to-analog converter first transforms the features into analog voltages. ACAM then performs in-memory search operations to retrieve the final result. The right part of Fig. 1 provides an overview of ACAM serving as an accelerator for the decision tree in the left part of Fig. 1. In contrast to TCAM, ACAM only requires three columns to represent the features, and each of the cells in ACAM stores an analog range instead.
However, despite its lower capacity requirement, ACAM demands additional peripheral circuits, and its larger cell size does not necessarily offer better area efficiency than TCAM. Furthermore, analog computing is susceptible to various non-idealities, such as stuck-at faults, IR drop, thermal noise, shot noise, and random telegraph noise [20]. While mitigation techniques exist [22], they often introduce additional memory overhead and offer limited effectiveness. Due to these reliability challenges, ACAM is not yet viable for real-world implementation. Therefore, this work opts for TCAM as the underlying hardware. Still, RETENTION remains compatible with ACAM after slight modifications (refer to Section III-E for detailed explanation).
@@ II-C Observation
To accelerate tree-based model inference using TCAM, each root-to-leaf path is mapped to a TCAM row, with each column representing a unique condition within the ensemble. Since a single path encounters only a minimal fraction of the conditions within the ensemble, the majority of TCAM cells store the X state for format alignment, leading to significant redundancy. In fact, mapping a tree-based model to TCAM without optimizations requires at least $\textit{paths}\times\#\textit{unique\_conditions}$ bits, which might exceed the memory capacity available in resource-constrained environments. As shown in Fig. 1, mapping a decision tree to a CAM with perfect capacity leads to over one-third of cells storing the X state. Moreover, due to the extremely high write latency of nvTCAM, real-time writing severely degrades performance, making it critical to fit the entire model into nvTCAM. Table I presents an experiment that highlights the substantial memory requirement and the overwhelming redundancy caused by the widespread use of the X state in TCAM. In this experiment, Random Forest models with 100 decision trees require up to 700MB of memory when mapped with naive unified mapping (refer to Section III-D1 for detailed explanation), and all exhibited redundancy of more than 96%. These results highlight that directly deploying nvTCAM as a tree-based model accelerator is both infeasible and inefficient in resource-constrained environments.
Previous work DT2CAM [35] introduced a framework for mapping decision trees to TCAM. However, instead of addressing memory redundancy, it only proposed an energy-saving precharge mechanism that disables mismatched rows to conserve power. Pedretti et al. [31] proposed mapping Random Forests to ACAM, incorporating techniques such as (1) limiting maximum tree depth and (2) reordering features while eliminating empty rows to mitigate memory redundancy. However, pre-pruning algorithms can result in severe accuracy degradation when implementing on bagging-based models (refer to Section III-B for detailed explanation), and alternative pruning strategies could achieve better accuracy within the same memory budget. Furthermore, although feature reordering groups redundant cells and empty-row elimination reduces capacity requirement, the associated computational overhead for merging partial match results is relatively high. In fact, comparable computational overhead could yield even greater memory savings. A follow-up study by Pedretti et al. [32] proposed an ACAM-based analog-digital architecture for tree-based ensemble model acceleration, assigning individual trees to separate ACAMs for architectural and compilation simplicity. While this approach effectively reduces redundancy across different trees, redundancy within each separate tree remains, leaving room for further optimization. Addressing this residual redundancy could lead to more reduction in memory consumption while maintaining computational efficiency.